OpenStudy (anikate):

angelina makes 70% of her free throws. What is the probability that she will NOT make make her nest two free throws?

OpenStudy (anikate):

i am going to wait

OpenStudy (anonymous):

.70 = probability of making a free throw 1 - .70 = probability of missing a free throw (1 - .70) ^2 = probability of missing two free throws (1 - .70)^2 = .3^2 = .09 9% chance

OpenStudy (anikate):

help me with this?

OpenStudy (anikate):

2 free throws?

OpenStudy (anonymous):

"What is the probability that she will NOT make make her nest TWO free throws?" With emphasis added

OpenStudy (anikate):

can we first do my regular question plz?

OpenStudy (anikate):

hold on 2 minutes

OpenStudy (anikate):

i will be back

OpenStudy (anikate):

nvm i got it cya